OpenStudy (anonymous):

simplify: 1/x+2 over 1+ 1/x+2

Nnesha (nnesha):

\[\huge\rm \frac{ \frac{ 1 }{ x+2 } }{ \frac{ 1+1 }{ x+2 }}\] like this ??

Nnesha (nnesha):

@BurntReynolds

OpenStudy (anonymous):

no